can someone help me solve 2^2x+1=15 for x
Answer by Alan3354(69443)

can someone help me solve 2^2x+1=15 for x
----------
2^2x+1=15
2^2x = 14
2x*log(2) = log(14)
x = log(14)/log(4)
x =~ 1.9037